please explain how you got your answers so I can understand
Answer #1

Ans 1) C ( linear momentum is conseved until no external force is applied)

Ans 2) C ( linear momentum conservervation ( m V (initial) = m1 V1 + m2 V2  (final) ) V is velocity vector

Ans 3) D ( imagine the disc and try to move fingers of your right hand in the direction of arrow in the diagram, The direction of torque as well as angular acceleration is the direction in which your thumb points, Right hand rule)

Ans 4)C ( using = v^{^{2}} = u^{_{2}} + 2 \alpha \theta , \theta = 2 \pi r , u =0 )

Ans 5) A ( torque = I * alpha, alphs from previous question, and I ( moment of inertia) = m * r2 (point mass at distance r )

Ans 6) D (torque = F x r; but here we are asked about the reaction at origin thus = - ( F x r))

Ans 7) B (Energy is conserved)

Ans 8) B ( Ring has highest moment of inertia, it gains the most amount of energy)